# Almost commuting self-adjoint operators and measurements

### v2

@@ 4 Multiple self-adjoint operators
Let be a unital -algebraβ and be an ideal of which is a -unital -algebraβ of real rank zero. Suppose that is a projection such that where is the quotient map. Then there is a projection such that
@@
Let be given. and Choose Since is a -near-spectrum of there is a unital c.p.c. map such that
| (eβ4.130) |
Since is a -spectrum of there exists a unital monomorphism such that
| (eβ4.131) |
@@
| (eβ4.132) |
It follows that in Let be a -dense subset of Let be mutually orthogonal non-zero projections in such that in Define by
| (eβ4.133) |
@@
| (eβ4.149) | |||||
Hence, using the inequality above and (eβ4.130),
| (eβ4.150) |
@@
It follows from Theorem 1.1 of [39] that and Therefore, (2), (3), (5), (6), (8) and (9) hold. β
@@
Let be a unital -algebra. Let and be maps, and We say satisfies condition (R, r, b, T, s) if has --cancellation with for all -stable rank -divisible rank and for all
@@
Let where is a compact subset. Let be maps and
For any there exist a positive number and an integer satisfying the following: Suppose that is a unital -algebraβ satisfying condition such that for any compact metric space and are unital two c.p.c. maps such that
| (eβ4.151) |
@@
Let and Since we assume that for any compact metric space for any finite subset as in Theorem 3.1 of [10], we have that for any sufficiently almost multiplicative c.p.c. maps Then, by Proposition 4.3 and applying Theorem 3.1 of [10], we obtain and a finite subset such that, if
| (eβ4.153) |
there exists an integer a unital homomorphismβ and a unitary such that
| (eβ4.154) |
@@
Let and Let and be maps and There exists satisfying the following: Suppose that is a unital -algebraβ such that for any compact metric space () satisfying condition and is an essential ideal which has real rank zero such that is purely infinite simple, for any projection and suppose that with () such that
| (eβ4.156) |
@@
| (eβ4.159) |
By the choice of (and ), by applying Proposition 2.11, and are not empty set (see also (eβ2.15)) and the -tuple has an -near-spectrum and, by Theorem 3.9, has an -spectrum Moreover (by the choice of ) such that (by applying Proposition 2.15),
| (eβ4.160) |
@@
By the choice of and (eβ4.162), as and applying Theorem 4.5, we obtain a unital homomorphismβ with finite dimensional range and a unitary such that
| (eβ4.163) |
@@
By the assumption that and (eβ4.160), is -dense in Note that
| (eβ4.166) |
@@
| (eβ4.175) |
Then, since by (eβ4.174),
| (eβ4.176) |
Since is a homomorphismβ so is Hence
| (eβ4.177) |
β
@@
Let be a -unital purely infinite simple -algebraβ and be a countably generated Hilbert -module. Let Then, by Kasparovβs absorbing theorem (Theorem 2 of [15]), is an orthogonal summand of and hence is isomorphic to a hereditary -subalgebraβ of which is purely infinite simple. If is unital, by Theorem 1 of [15], Then the conclusion of Theorem 1.2 follows from Theorem 3.9 (without referring to synthetic spectra).
Otherwise, since is countably generated, is non-unital but -unital (see Proposition 3.2 of [2]). By Theorem 1 of [15], Note that (see Remark 2.5 (c) of [30]) which has real rank zero by [36]. It follows from Theorem 3.3 of [37] (see also Remark 2.5 (b) of [30]) that is purely infinite simple, and every projection in is equivalent to Thus, by Proposition 4.3, Theorem 4.6 applies. β
@@ 5 A pair of almost commuting self-adjoint operators on Hilbert modules
For a unital -algebraβ denote by the set of elements with the property that for no ideal of is one-sided and not two-sided invertible in Recall that is said to have property (IR) if all elements in belongs to the norm closure of the group of invertible elements.
A non-unital -algebraβ has property (IR), if the -algebraβ obtained by adjoining a unit to has property (IR).
@@
Every -algebraβ of stable rank one has property (IR), and every purely infinite simple -algebraβ has property (IR).
@@
(Note that the identity in (eβ5.179) also implies that )
@@
| (eβ5.180) |
there exists a compact subset and a unital injective homomorphismβ such that
| (eβ5.181) |
@@
| (eβ5.183) |
It follows that Hence and are in the same path connected component of It follows that β
@@
Let There exists satisfying the following: Suppose that is a unital -algebraβ with an essential ideal which is a -unital -algebraβ with real rank zero and with property (IR) such that is purely infinite simple. Suppose that with () such that
| (eβ5.184) |
@@
Recall that is a purely infinite simple -algebra. Therefore, by the choice of and applying Theorem 5.2 we obtain a compact subset a unital injective homomorphismβ such that
| (eβ5.186) |
@@
| (eβ5.188) |
To see the claim, we note that, by (eβ5.187), Then, by (eβ5.186) and by the assumption (eβ5.184),
| (eβ5.189) |
@@
By Corollary 3.7, there is a unital injective homomorphismβ such that
| (eβ5.191) |
@@
To see this, let us assume it is not the case (i). Then is in a bounded path connected open component ButοΌ since one must have Hence
@@
| (eβ5.193) |
Hence, by the second part of (eβ5.191),
| (eβ5.194) |
If and (i.e., in the case (ii)), by the second part of (eβ5.191) and the claim above, one computes that
| (eβ5.195) |
@@
| (eβ5.197) |
Then, by (eβ5.186), (eβ5.191) and (eβ5.197),
| (eβ5.198) | |||||
| (eβ5.199) |
We may write where and are mutually orthogonal non-zero projections in and There are two cases: (1) has real rank zero, then, by Theorem 3.1 of [3], every projection in lifts. If is path connected, by the βMoreoverβ part of Theorem 3.5, we may arrange such that or in Since is unital, Thus, by Lemma 4.1, in both cases, there are mutually orthogonal projections such that There exists such that
| (eβ5.200) |
@@
| (eβ5.202) |
Moreover, by the second part of (eβ5.201) and (eβ5.200), and then, by the first part of (eβ5.201), (eβ5.200), (eβ5.202), we obtain
| (eβ5.203) | |||
| (eβ5.204) | |||
| (eβ5.205) |
@@
| (eβ5.206) |
It follows from (eβ5.204) and (eβ5.205), and by the assumption on and that
| (eβ5.207) | |||
| (eβ5.208) |
@@
For (ii), let and If is unital, then As in 4.7, this case follows from Theorem 3.9 (regardless of Fredholm index). Also in the proof of Theorem 1.2 (see 4.7), when is not unital, it is a -unital purely infinite simple -algebraβ which has real rank zero and has property (IR). The fact that is purely infinite is also discussed in 4.7 (the proof of Theorem 1.2). Thus Theorem 5.3 applies.
For (iii), let and As in 4.7 (the proof of Theorem 1.2), by Kasparovβs theorems, and is a -unital hereditary -subalgebraβ of a -unital simple -algebraβ of real rank zero and stable rank one. Hence is also a simple -algebraβ of real rank zero and stable rank one, whence has property (IR). If is unital, then Therefore the conclusion follows from Theorem 4.4 of [8]. Otherwise, by the assumption, has continuous scale (see of [27]). Hence, by Corollary 3.3 of [27], is a simple purely infinite simple -algebra. Thus Theorem 5.3 applies. β
@@
The next proposition provides a short way to see the second condition in (eβ5.184) is necessary. We also provide a direct proof.
@@
| (eβ5.216) |
Note that (eβ5.216) also implies that is invertible.
@@
| (eβ5.218) |
where By (eβ2.16), we may further have
| (eβ5.219) |
@@
| (eβ5.220) |
Now if then, by (eβ5.219), Since is normal,
| (eβ5.221) |
Hence, by (eβ5.220),
| (eβ5.222) |
@@
On the other hand, computationally speaking, -synthetic-spectrum could be tested. Note that a -synthetic-spectrum consists of finitely many balls and there are only finitely many bounded connected components in its complement. This allows one to avoid the situation of infinitely many βholesβ in the spectrum. Note that the index in each of these βholesβ remains constant. So one only needs to test one point in each of these βholesβ. One also notes that the second condition in (eβ5.184) does not imply
@@
In Theorem 1.1 of [17], seems depending on a given operator The constant in Theorem 5.3 does not depend on nor even -algebras (as long as satisfies the conditions in Theorem 5.3). Finally, let us note that when a unital -algebraβ has (IR), then, by Theorem 4.4 of [8], a pair of almost commuting self-adjoint elements and is close to a pair of commuting ones even though may be large (see also the example below).
@@
Let be a unital separable purely infinite simple -algebraβ with and and be a non-unital separable simple -algebraβ with We assume that either is purely infinite simple, or is a separable simple -algebraβ of real rank zero and stable rank one. Consider any unital -algebraβ which is given by the following essential extension:
@@ 6 Epilogue
The next corollary shows that, when the second condition in (eβ4.156) is stronger than that of vanishing index in Theorem 1.3.
@@
We choose such that the conclusion of Proposition 2.15 holds for Then we choose such that the conclusion of Proposition 2.11 holds for Choose Suppose that (eβ6.225) holds. Put Thus (applying Proposition 2.11) and are all non-empty sets. By Proposition 2.15,
| (eβ6.227) |
@@
Applying Theorem 4.6, one obtains commuting -tuple of self-adjoint operators Let be the -subalgebraβ of generated by Define to be the embedding. Since for any compact metric space we conclude that Consequently where β
@@
| (eβ6.246) |
then one obtains a sequence of homomorphisms such that
| (eβ6.247) |
@@
| (eβ6.252) |
Let be the -subalgebraβ of generated by Then there is an isomorphism for some compact subset such that and Define by for Since generates we have
| (eβ6.253) |
@@
| (eβ6.256) |
This contradicts (eβ6.238). Therefore there is no homomorphisms such that (eβ6.248). In other words, there is such that
| (eβ6.257) |
where the infimum is taken among all homomorphisms Combining this with (eβ6.245), the proposition follows. β
@@
In quantum theory, one may view macroscopic observables as some self-adjoint operators on a Hilbert space which we assume to be infinite dimensional and separable. Let be the unital -subalgebraβ generated by and (let us assume for convenience here). Suppose that is an orthonormal basis of and suppose that, for any there is such that for all In this case these observables collapse uniformly and any measurements outside (orthogonally) a finite dimensional subspace vanish. In other words, outsiders would see nothing. Let us assume that none of these observables collapse this way (see also Remark 6.5). In other words, we assume that none of βs are compact.
Given a state which is defined by a unit vector the expected value when measuring an observable given by an operator is (see p.178 of [33]). βApproximately Macroscopically Uniqueβ states introduced by Mumford (see Chapter 14 of [33]), or AMU states, are βnear eigenvectorsβ. Set, for a small
where As pointed out in [33], for AMU to be non-empty, it is natural to assume that the commutators have small bound. If the question in the introduction (see (eβ1.1) and (eβ1.2)) had an affirmative answer, then the quantum system could be approximated by a classical one. Consequently, AMU does exist. Observed by Mumford, the condition that the commutators are small is not quite enough (Proposition 6.4 further shows the complexity of the problem) in general.
In an infinite dimensional separable Hilbert space a self-adjoint operator such as may not have any eigenvalues (but have approximate eigenvalues). However, if one allows non-normal states of or that of then does have eigenstates and eigenvalues in the atomic representation of or (see, for example, Proposition 4.3.10 of [34]). In particular, one should consider those states of which vanish on the ideal of compact operators. If a point in the unit cube of is expected to be a βjoint eigenvalueβ, then one may use the test function defined in 2.7 to test it. In order to have (for small ), one only needs one vector state (with a unit vector ) such that
| (eβ6.258) |
One may expect this can be repeatedly measured. Suppose that are unit vectors and is a finite dimensional subspace spanned by Suppose that there is another unit vector such that (eβ6.258) also holds for (in place of ), and this persists. Then we obtain a sequence of () such that (eβ6.258) holds for each (in place of ). Note that unit vector states given by has limit points which give states of vanishing on and Note that there are still unknown (approximate) eigenvalues outside any finite dimensional subspaces of
@@
| (eβ6.260) |
If whenever (eβ6.258) holds, inequality (eβ6.259) holds for some quantum states then the second condition in (eβ1.3) always satisfies (in fact, in this case ).
Let us, for a moment, call a unit vector state for which (eβ6.258) holds a synthetic-AMU state and a synthetic-joint-expected value. The state for which (eβ6.259) holds may be called an essential synthetic-AMU and (for which (eβ6.259) holds) an essential synthetic-joint-expected value. Note that it is easy to see an AMU state (with sufficiently small ) is an ()-synthetic-AMU state. Moreover, by Proposition 2.11, the set of synthetic-AMU states is non-empty, whenever is small (independent of ). The second condition in Theorem 1.1 then may be stated as follows:

### v3

@@ 4 Multiple self-adjoint operators
Let be a unital -algebra and be an ideal of which is a -unital -algebra of real rank zero. Suppose that is a projection such that where is the quotient map. Then there is a projection such that
@@
Let Choose Since is a -near-spectrum of there is a unital c.p.c. map such that
| (e 4.130) |
Since is a -spectrum of there exists, by Corollary 3.7, a compact subset such that and a unital monomorphism such that
| (e 4.131) |
@@
| (e 4.132) |
It follows that in Let be a -dense subset of which is also -dense in Let be mutually orthogonal non-zero projections in such that in Define by
| (e 4.133) |
@@
| (e 4.148) | |||||
| (e 4.149) |
Hence, using the inequality above and (e 4.130),
| (e 4.150) |
@@
It follows from Theorem 1.1 of [40] that and Therefore, (2), (3), (5), (6), (8) and (9) hold. ∎
@@
Let be a unital -algebra. Let and be maps, and We say satisfies condition if has --cancellation with for all -stable rank -divisible rank for all and for all
@@
Let where is a compact subset. Let be maps, and
For any there exist a positive number and an integer satisfying the following: Suppose that is a unital -algebra satisfying condition such that for any compact metric space and are unital c.p.c. maps such that
| (e 4.151) |
@@
Let and Since we assume that for any compact metric space for any finite subset as in Theorem 3.1 of [10], we have that for any sufficiently multiplicative c.p.c. maps Then, by Proposition 4.3 and applying Theorem 3.1 of [10], we obtain and a finite subset such that, if
| (e 4.153) |
there exists an integer a unital homomorphism with finite dimensional range and a unitary such that
| (e 4.154) |
@@
Let and Let and be maps and There exists satisfying the following: Suppose that is a unital -algebra such that for any compact metric space () satisfying condition and is an essential ideal which has real rank zero such that is purely infinite simple, for any non-zero projection and suppose that with () such that
| (e 4.156) |
@@
| (e 4.159) |
By the choice of (and ) and, applying Proposition 2.11, and are non-empty set (see also (e 2.15)) and the -tuple has an -near-spectrum and, by Theorem 3.9, has an -spectrum Moreover (by the choice of ), we have (by applying Proposition 2.15),
| (e 4.160) |
@@
By the choice of and (e 4.162), as and applying Theorem 4.5, we obtain a unital homomorphism with finite dimensional range and a unitary such that
| (e 4.163) |
@@
By the assumption that and (e 4.160), is -dense in Note that
| (e 4.166) |
@@
| (e 4.176) |
Then, since by (e 4.175),
| (e 4.177) |
Since is a homomorphism so is Hence
| (e 4.178) |
@@
Let be a -unital purely infinite simple -algebra and be a countably generated Hilbert -module. Let Then, by Kasparov’s absorbing theorem (Theorem 2 of [15]), is an orthogonal summand of and hence is isomorphic to a hereditary -subalgebra of which is purely infinite simple. If is unital, by Theorem 1 of [15], Then the conclusion of Theorem 1.2 follows from Theorem 3.9 (without referring to synthetic spectra).
Otherwise, since is countably generated, is non-unital but -unital (see Proposition 3.2 of [2]). By Theorem 1 of [15], Note that (see Remark 2.5 (c) of [31]) which has real rank zero by [37]. It follows from Theorem 3.3 of [38] (see also Remark 2.5 (b) of [31]) that is purely infinite simple, and every projection in is equivalent to Thus, by Proposition 4.3, Theorem 4.6 applies. ∎
@@ 5 A pair of almost commuting self-adjoint operators on Hilbert modules
For a unital -algebra denote by the set of elements with the property that for no ideal of is one-sided and not two-sided invertible in Recall that is said to have property (IR) if all elements in belongs to the norm closure of the group of invertible elements.
A non-unital -algebra has property (IR), if the -algebra obtained by adjoining a unit to has property (IR).
@@
Every -algebra of stable rank one has property (IR), and every purely infinite simple -algebra has property (IR).
@@
(Note that the identity in (e 5.180) also implies that )
@@
| (e 5.181) |
there exists a compact subset and a unital injective homomorphism such that
| (e 5.182) |
@@
| (e 5.184) |
It follows that Hence and are in the same path connected component of It follows that ∎
@@
Let There exists satisfying the following: Suppose that is a unital -algebra with an essential ideal which is a -unital -algebra with real rank zero and with property (IR) such that is purely infinite simple. Suppose that with () such that
| (e 5.185) |
@@
Recall that is a purely infinite simple -algebra. Therefore, by the choice of and applying Theorem 5.2 we obtain a compact subset a unital injective homomorphism such that
| (e 5.187) |
@@
| (e 5.189) |
To see the claim, we note that, by (e 5.188), Then, by (e 5.187) and by the assumption (e 5.185),
| (e 5.190) |
@@
By Corollary 3.7, there is a unital injective homomorphism such that
| (e 5.192) |
@@
To see this, let us assume it is not the case (i). Then is in a bounded path connected open component But, since one must have Hence
@@
| (e 5.194) |
Hence, by the second part of (e 5.192),
| (e 5.195) |
If and (i.e., in the case (ii)), by the second part of (e 5.192) and the claim above, one computes that
| (e 5.196) |
@@
| (e 5.198) |
Then, by (e 5.187), (e 5.192) and (e 5.198),
| (e 5.199) | |||||
| (e 5.200) |
We may write where and are mutually orthogonal non-zero projections in and There are two cases: (1) has real rank zero, then, by Theorem 3.14 of [3], every projection in lifts. (2) If is path connected, by the “Moreover” part of Theorem 3.5, we may arrange such that or in Since is unital, Thus, by Lemma 4.1, in both cases, there are mutually orthogonal projections such that There exists such that
| (e 5.201) |
@@
| (e 5.203) |
Moreover, by the second part of (e 5.202) and (e 5.201), and then, by the first part of (e 5.202), (e 5.201), (e 5.203), we obtain
| (e 5.204) | |||
| (e 5.205) | |||
| (e 5.206) |
@@
| (e 5.207) |
It follows from (e 5.205) and (e 5.206), and by the assumption on and that
| (e 5.208) | |||
| (e 5.209) |
@@
For (ii), let and If is unital, then As in 4.7, this case follows from Theorem 3.9 (regardless of Fredholm index). Also in the proof of Theorem 1.2 (see 4.7), when is not unital, it is a -unital purely infinite simple -algebra which has real rank zero and has property (IR). The fact that is purely infinite is also discussed in 4.7 (the proof of Theorem 1.2). Thus Theorem 5.3 applies.
For (iii), let and As in 4.7 (the proof of Theorem 1.2), by Kasparov’s theorems, and is a -unital hereditary -subalgebra of a -unital simple -algebra of real rank zero and stable rank one. Hence is also a simple -algebra of real rank zero and stable rank one, whence has property (IR). If is unital, then Therefore the conclusion follows from Theorem 4.4 of [8]. Otherwise, by the assumption, has continuous scale (see of [27]). Hence, by Corollary 3.3 of [27], is a simple purely infinite simple -algebra. Thus Theorem 5.3 applies. ∎
@@
The next proposition provides a short way to see the second condition in (e 5.185) is necessary. We also provide a direct proof.
@@
| (e 5.217) |
Note that (e 5.217) also implies that is invertible.
@@
| (e 5.219) |
where By (e 2.16), we may further have
| (e 5.220) |
@@
| (e 5.221) |
Now if then, by (e 5.220), Since is normal,
| (e 5.222) |
Hence, by (e 5.221),
| (e 5.223) |
@@
On the other hand, computationally speaking, -synthetic-spectrum could be tested. Note that a -synthetic-spectrum consists of finitely many balls and there are only finitely many bounded connected components in its complement. This allows one to avoid the situation of infinitely many “holes” in the spectrum. Note that the index in each of these “holes” remains constant. So one only needs to test one point in each of these “holes”. One also notes that the second condition in (e 5.185) does not imply
@@
In Theorem 1.1 of [17], seems depending on a given operator The constant in Theorem 5.3 does not depend on nor even -algebras (as long as satisfies the conditions in Theorem 5.3). Finally, let us note that when a unital -algebra has (IR), then, by Theorem 4.4 of [8], a pair of almost commuting self-adjoint elements and is close to a pair of commuting ones even though may be large (see also the example below).
@@
Let be a unital separable purely infinite simple -algebra with and and be a non-unital separable simple -algebra with We assume that either is purely infinite simple, or is a separable simple -algebra of real rank zero and stable rank one. Consider any unital -algebra which is given by the following essential extension:
@@ 6 Epilogue
The next corollary shows that, when the second condition in (e 4.156) is stronger than that of vanishing index in Theorem 1.3.
@@
We choose such that the conclusion of Proposition 2.15 holds for Then we choose such that the conclusion of Proposition 2.11 holds for Choose Suppose that (e 6.226) holds. Put Thus (applying Proposition 2.11) and are all non-empty sets. By Proposition 2.15,
| (e 6.228) |
@@
Applying Theorem 4.6, one obtains commuting -tuple of self-adjoint operators Let be the -subalgebra of generated by Define to be the embedding. Since for any compact metric space we conclude that Consequently where ∎
@@
| (e 6.247) |
then one obtains a sequence of unital homomorphisms such that
| (e 6.248) |
@@
| (e 6.253) |
Let be the (commutative) -subalgebra of generated by Then there is an isomorphism for some compact subset such that and Define by for Since generates we have
| (e 6.254) |
@@
| (e 6.257) |
This contradicts (e 6.239). Therefore there is no homomorphisms such that (e 6.249). In other words, there is such that
| (e 6.258) |
where the infimum is taken among all homomorphisms Combining this with (e 6.246), the proposition follows. ∎
@@
In quantum theory, one may view macroscopic observables as some self-adjoint operators on a Hilbert space which we assume to be infinite dimensional and separable. Let be the unital -subalgebra generated by and (let us assume for convenience here). Suppose that is an orthonormal basis of and suppose that, for any there is such that for all In this case these observables collapse uniformly and any measurements outside (orthogonally) a finite dimensional subspace vanish. In other words, outsiders would see nothing. Let us assume that none of these observables collapse this way (see also Remark 6.5). In other words, we assume that none of ’s are compact.
Given a state which is defined by a unit vector the expected value when measuring an observable given by an operator is (see p.178 of [34]). “Approximately Macroscopically Unique” states introduced by Mumford (see Chapter 14 of [34]), or AMU states, are “near eigenvectors”. Set, for a small
where As pointed out in [34], for AMU to be non-empty, it is natural to assume that the commutators have small bound. If the question in the introduction (see (e 1.1) and (e 1.2)) had an affirmative answer, then the quantum system could be approximated by a classical one. Consequently, AMU does exist. Observed by Mumford, the condition that the commutators are small is not quite enough (Proposition 6.4 further shows the complexity of the problem) in general.
In an infinite dimensional separable Hilbert space a self-adjoint operator such as may not have any eigenvalues (but have approximate eigenvalues). However, if one allows non-normal states of or that of then does have eigenstates and eigenvalues in the atomic representation of or (see, for example, Proposition 4.3.10 of [35]). In particular, one should consider those states of which vanish on the ideal of compact operators. If a point in the unit cube of is expected to be a “joint eigenvalue”, then one may use the test function defined in 2.7 to test it. In order to have (for small ), one only needs one vector state (with a unit vector ) such that
| (e 6.259) |
One may expect this can be repeatedly measured. Suppose that are unit vectors and is a finite dimensional subspace spanned by Suppose that there is another unit vector such that (e 6.259) also holds for (in place of ), and this persists. Then we obtain a sequence of () such that (e 6.259) holds for each (in place of ). Note that unit vector states given by has limit points which give states of vanishing on and Note that there are still unknown (approximate) eigenvalues outside any finite dimensional subspaces of
@@
| (e 6.261) |
If whenever (e 6.259) holds, inequality (e 6.260) holds for some quantum states then the second condition in (e 1.3) always satisfies (in fact, in this case ).
Let us, for a moment, call a unit vector state for which (e 6.259) holds a synthetic-AMU state and a synthetic-joint-expected value. The state for which (e 6.260) holds may be called an essential synthetic-AMU and (for which (e 6.260) holds) an essential synthetic-joint-expected value. Note that it is easy to see an AMU state (with sufficiently small ) is an ()-synthetic-AMU state. Moreover, by Proposition 2.11, the set of synthetic-AMU states is non-empty, whenever is small (independent of ). The second condition in Theorem 1.1 then may be stated as follows: